OpenStudy (anonymous):

two fair dice are rolled, what is the probability that the total showing is either even or less than four?

OpenStudy (tommynaut):

Have a go yourself first. Do you understand what "and" and "or" mean in probability questions? Usually, "and" implies multiplication (+), "or" implies addition (+). So, for example, if we have "what is the probability of rolling a 1 OR a 2 on a fair dice?", we calculate the probability of rolling a 1 and we add that to the probability of rolling a 2, because either are allowed. So, that would be 1/6 + 1/6 = 2/6 = 1/3. Try and use a similar sort of method for your problem. Work out the probability of the two rolled dice showing an even number, then add that to the probability of the dice values added together being less than 4.